OpenStudy (kpop4life123):

A ball bearing is shaped like a sphere and has a diameter of 2.5 centimetres. What is the volume contained inside the ball bearing? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 8.18 cubic centimeters 7.15 cubic centimeters 7.08 cubic centimeters 6.89 cubic centimeters

OpenStudy (kendricklamar2014):

\[\frac{ 2.5 }{ 2 } = ?\] @Kpop4life123

OpenStudy (kpop4life123):

1.25?

OpenStudy (kendricklamar2014):

That is correct...

OpenStudy (kendricklamar2014):

Volume of a Sphere formula: \[V = \frac{ 4 }{ 3 } \pi r ^3\]

OpenStudy (kpop4life123):

i dont actually know how to do this?

OpenStudy (kendricklamar2014):

Now just substitute: \[V = \frac{ 4 }{ 3 } \times 3.14 \times 1.25^3\]

OpenStudy (kpop4life123):

oh ok wait let me try it

OpenStudy (kendricklamar2014):

Ok

OpenStudy (kpop4life123):

8.18?

OpenStudy (kendricklamar2014):

That is correct

OpenStudy (kpop4life123):

thanks ! :)

OpenStudy (kendricklamar2014):

No problem =)